# AzSLD: Azerbaijani Sign Language Dataset for Fingerspelling, Word, and Sentence Translation with Baseline Software

### v1

@@ 3.1 Dataset components
AzSLD Words comprises 100 classes of word videos, representing the top 100 most frequent words observed within the sentences of the AzSLD Sentences dataset. The number of samples in the word-level dataset is highly imbalanced across classes. Therefore, we recommend using class weights for accuracy measurement in word-level recognition models. The recording setup for word-level videos is the same as that used for sentence-level videos.
@@
The strategy used for the fingerspelling dataset is not applicable to the collection of the sentence dataset; the interpretation of each sentence and demonstration of the signs need to be moderated. AzSLD Sentences comprises signing videos for 500 sentences, collected using frontal and side-view cameras. Each dataset folder contains time-aligned sentence annotation files and two subfolders for the recordings from the two cameras. The videos were performed by 18 to 25 different signers, with a slight imbalance among them.
A custom setup was used for filming, with a white background and a fixed seating distance from the camera. To obtain more information from the collected data, it was decided to capture the video using two cameras: one mounted on the monitor, looking straight at the actor, and another mounted on the top-left side. These two views are considered complimentary - those signs that are not visible from the front view could be noticed in the side view and vice versa (e.g., which finger has been folded or what the mouth mactivity when a hand accidentally covers it). Fig. 1 demonstrates examples of screens from both cameras for 4 different signs. Video files from the two cameras have identical names and are located in the ’Cam1’ and ’Cam2’ folders, respectively in the sentence-level dataset.
@@
The directory structure of the AzSLD is as follows:
[…]
.1 AzSLD_Fingerspelling/. .2 A/. .3 a1.jpg. .3 …. .3 a450.jpg. .2 B/. .2 C/. .2 …. .2 Z/.
[…]
.1 AzSLD_Sentence/. .2 1/. .3 ann/. .4 2022-06-16 15-44-27.mp4.json. .4 …. .4 2022-07-18 12-21-14.mp4.json. .3 Cam1/. .4 2022-06-16 15-44-27.mp4. .4 …. .4 2022-07-18 12-21-14.mp4. .3 Cam2/. .4 2022-06-16 15-44-27.mp4. .4 …. .4 2022-07-18 12-21-14.mp4. .2 …. .2 500.
@@ 3.2 Context, linguistic analysis, and annotations in AzSLD Sentences
To enhance usability and accessibility, the sentences were simplified. By simplifying the language used in the dataset, it becomes more inclusive and caters to a wider range of users, regardless of their sign language proficiency level. Each sentence was recorded by at least 18 of the 40 participants involved in the project. The involvement of multiple signers ensured that the dataset captured a wide range of signing styles, signing speed variations, and nuances. This approach also accounts for individual differences in signing preferences and regional variations within Azerbaijani Sign Language.
Unlike spoken languages, sign languages lack computational tools for performing semi-automatic or automatic lexical annotations [57]. This makes the annotation of sign language (SL) datasets a labor-intensive manual task. In this study, we present the AzSL dataset, developed in alignment with recently established agreements on annotation standards in recent research.
Sign languages have specific linguistic rules that do not directly reflect spoken language [58]. Similar to the problem of translation between spoken languages, which opposes two sequences of variable length and tokens, sign language translation requires an understanding of the meaning and order of signs in a video [59]. In an isolated sentence-level annotation, each distinct label, such as a word, number, or letter, is annotated with respect to the corresponding frame number in the video. This type of annotation provides a broader context for feature extraction and allows the analysis of sentence-level features and structures. The total length of the video varied according to the number of words performed in the video. 2 visualizes the variability of the length of videos in frame numbers.
@@ 3.4 Data loader for ease of training and testing
Data loaders are a fundamental component in machine learning pipelines, serving as the bridge between raw data and model training. The lack of well-functioning data loaders hinders their usability for the machine learning community [64]. Providing a data loader for new sign language datasets is essential for the research community, particularly for benchmarking purposes. We have developed a data loader for AzSLD that can be adapted for other sign language datasets as well. By addressing the unique challenges posed by the dataset and adhering to best practices, the provided data loader can significantly enhance the efficiency and effectiveness of machine learning models in this domain. When developing methods for data retrieval and traversal, the following factors should be considered:
  • Speed of the data retrieval. It should be possible to utilize the possibilities, features and capacity of the programming language, environment, and hardware to achieve the highest access to the data.
  • Easiness of use by the training models. The data loading and retrieval processes shall consider the easy integration with the training environment. This decision shall not merely consider the programming language or frameworks (Tensorflow or Pytorch) but also the domain of the application. It is needless to consider lightweight ML frameworks if the target is complex spatio-temporal data processing.
  • Customization of the details. The environment of the end user will be different from the one that dataset developer has used. To free the user from digging and understanding the principle of data loader’s work, it is important to design the program based on configurable values and actions. It should also be possible to include or exclude the particular processing steps and require the output in one of the suggested formats. None of the model parameters, like batch size or layers should be hardcoded.
  • Openness of the system. The loading and traversal system shall not be introduced as an external library to the users. It shall be easy to include, embed and customize the code to tailor it for the required purpose.

### v2

@@ 3.1 Dataset components
AzSLD Words comprises 100 classes of word videos, representing the top 100 most frequent words observed within the sentences of the AzSLD Sentences dataset. The number of samples in the word-level dataset is highly imbalanced across classes. Therefore, we recommend using class weights for accuracy measurement in word-level recognition models. The recording setup for word-level videos is the same as that used for sentence-level videos. Folder names indicate the ground truth labels for the ease of word-level model evaluation.
@@
The strategy used for the fingerspelling dataset is not applicable to the collection of the sentence dataset; the interpretation of each sentence and demonstration of the signs need to be moderated. AzSLD Sentences comprises signing videos for 500 sentences, collected using frontal and side-view cameras. Each dataset folder contains time-aligned sentence annotation files and two subfolders for the recordings from the two cameras. Ground truth annotations of sentences for each class were added in a separate file. The videos were performed by 18 to 25 different signers, with a slight imbalance among them.
A custom setup was used for filming, with a white background and a fixed seating distance from the camera. RGB videos were collected in consideration of the results for human-action recognition studies with depth cameras [55]. This type of setup makes a seamless application of transfer learning if a model has been pre-trained with an RGB or depth videos of a human-action dataset.
@@
To obtain more information from the collected data, it was decided to capture the video using two cameras: one mounted on the monitor, looking straight at the actor, and another mounted on the top-left side. These two views are considered complimentary - those signs that are not visible from the front view could be noticed in the side view and vice versa (e.g., which finger has been folded or what the mouth mactivity when a hand accidentally covers it). Fig. 1 demonstrates examples of screens from both cameras for 4 different signs. Video files from the two cameras have identical names and are located in the ’Cam1’ and ’Cam2’ folders, respectively in the sentence-level dataset.
@@ 3.2 Context, linguistic analysis, and annotations in AzSLD Sentences
To enhance usability and accessibility, the sentences were simplified. Simplifying the language used in the dataset makes it more inclusive and caters to a wider range of users, regardless of their sign language proficiency level. Each sentence was recorded by at least 18 of the 40 participants involved in the project. The involvement of multiple signers ensured that the dataset captured a wide range of signing styles, signing speed variations, and nuances. This approach also accounts for individual differences in signing preferences and regional variations within Azerbaijani Sign Language.
Sign languages have specific linguistic rules that do not directly reflect spoken language [58, 59]. Similar to the problem of translation between spoken languages, which opposes two sequences of variable length and tokens, sign language translation requires an understanding of the meaning and order of signs in a video [60]. In an isolated sentence-level annotation, each distinct label, such as a word, number, or letter, is annotated with respect to the corresponding frame number in the video. This annotation type provides a broader context for feature extraction and allows the analysis of sentence-level features and structures. The total length of the video varied according to the number of words performed in the video. 2 visualizes the variability of the length of videos in frame numbers.
@@ 3.4 Data loader for ease of training and testing
Data loaders are a fundamental component in machine learning pipelines, serving as the bridge between raw data and model training. The lack of well-functioning data loaders hinders their usability for the machine learning community [65]. Providing a data loader for new sign language datasets is essential for the research community, particularly for benchmarking purposes. We have investigated existing datasets as [66, 67] and adopted all the best in them. We have developed a data loader for AzSLD that can be adapted for other sign language datasets as well. By addressing the unique challenges posed by the dataset and adhering to best practices, the provided data loader can significantly enhance the efficiency and effectiveness of machine learning models in this domain. This data loader ensures the uniformity of input data, enabling seamless integration with machine learning pipelines.
Frame Extraction and Preprocessing
The data loader extracts a fixed number of frames ($n_{\text{frames}}$) from each video file using a specified temporal resolution controlled by a frame step parameter. The frames are resized and padded to a uniform resolution ($224\times 224$ pixels by default) while retaining the central region of interest. To ensure consistency across the dataset, the extracted frames are normalized and converted to a float32 format. The frame extraction process handles videos of varying lengths by dynamically adjusting the starting point and padding with zero frames if required.
Label Encoding
The loader supports label encoding by dynamically mapping categorical labels to one-hot encoded vectors. This is achieved through the creation of two dictionaries:
  • Label-to-One-Hot Dictionary: Maps each label to its corresponding one-hot encoded vector.
  • Index-to-Label Dictionary: Provides a reverse mapping from the index of the one-hot vector to the original label.
These mappings are saved and loaded using the pickle library, enabling reusability across different runs.
Dataset Splitting
The data loader incorporates a splitting function that divides the dataset into training and testing subsets. By default, 80% of the data is allocated to the training set, with the remaining 20% reserved for testing. This ensures a standardized pipeline for model evaluation and performance comparison.
Integration and Robustness
The data loader is designed to handle diverse datasets with varying directory structures and video properties. Labels are extracted from directory names, enabling compatibility with hierarchically organized datasets. Furthermore, the implementation is robust to inconsistencies, such as videos with insufficient frames, ensuring that all data is formatted correctly for model training.
Benefits and Applications
This data loader simplifies the preprocessing of sign language video datasets, making it suitable for various tasks such as gesture classification, action recognition, and translation systems. It ensures uniform data preparation, enhances reproducibility, and reduces the manual effort required to handle video data. This tool is a critical component of our pipeline, contributing to the efficient and accurate training of models for sign language recognition.